OpenStudy (anonymous):

initial value problem: 2y' * sqroot(x) = y y(0) = 1

OpenStudy (anonymous):

\[2y'\sqrt{x} = y\] y(0) = 1

OpenStudy (anonymous):

\[2dy/dx \sqrt{x}=y,\] dy/y=dx/(2(sqrt(x), \[\ln \left| y \right|=\sqrt{x}+c\] subst. (0,1) to get ln(1)=sqrt(0) +c so c=0

OpenStudy (anonymous):

Im hoping you know how to integrate , good luck!

OpenStudy (anonymous):

thank you lbt =3